Solve each system by substitution.
y=-3x+9
-4x+2y=-2

Answer by ewatrrr(24785)


Hi,
Solving the system by substitution
y = -3x+9 green line
-4x+2y=-2 blue line
-4x+2(-3x+9)= -2
-4x -6x + 18 = -2
-10x = -20
x = 2
y = -3x+9
y = -6+9
y = 3